help wih these question please3. Consider the following autoregressive process Yt = Bo + B1yt-1 + B2Yt-2 + Ut, where E (UtYt-1, Yt-2, ...) = 0. You obtaine
Answer #1

3) Since the estimate of the parameters are given the estimated regression model is:

Ey)= y = -0.2 + 0.4 % yt-1- 0.1 % yt-2

Now it is given that y19 = – 0.2 and y20 = 0.3

Therefore, EY21) = -0.2 + 0.4 * Y20 - 0.1 * 719

         Or, Yz1 = -0.2 +0.4*0.3 -0.1 *(-0.2) = -0.06

Thus,     EY22) = -0.2 + 0.4 * Y21 -0.1 * 720

        Or, EY22) = -0.2 +0.4*(-0.06) -0.1*(0.3) = -0.254

So, option (e) - 0.254 is the correct one.

4) The estimated regression model is:

             gr, = 27.34 + 2.97 * sthi - 2.04 * awbc;

Since Jimmy studies on average 11 hours per week so for Jimmy sth = 11 and he drinks on average 4.3 pints of beer a week so awbc = 4.3

Therefore, Jimmy’s predicted grade in ECON20110 is:

  gr = 27.34 + 2.97 * 11 - 2.04 * 4.3 = 51.24

So, option (e) 51.24 is the correct one.